# Communication-Efficient Distributed Deep Learning via Federated Dynamic Averaging

### v1

@@ 4.1. Setup
Platform. We employ TensorFlow (Abadi et al., 2015), integrated with Keras (Chollet et al., 2015), as the platform for conducting our experiments. We modified TensorFlow to include our FDA variants and all competitive algorithms.
@@
Datasets & Models. The core experiments involve training Convolutional Neural Networks (CNNs) of varying sizes and complexities on two datasets: MNIST (Deng, 2012) and CIFAR-10 (Krizhevsky, 2012). For the MNIST dataset, we employ LeNet-5 (Lecun et al., 1998), composed of approximately 62 thousand parameters, and a modified version of VGG16 (Simonyan and Zisserman, 2014), denoted as VGG16*, consisting of 2.6 million parameters. VGG16* was specifically adapted for the MNIST dataset, a less demanding learning problem compared to ImageNet (Russakovsky et al., 2015), for which VGG16 was designed. In VGG16*, we omitted the $512$-channel convolutional blocks and downscaled the final two fully connected (FC) layers from $4096$ to $512$ units each. Both models use Glorot uniform initialization (Glorot and Bengio, 2010). For CIFAR-10, we utilize DenseNet121 and DenseNet201 (Huang et al., 2016), as implemented in Keras (Chollet et al., 2015), with the addition of dropout regularization layers at rate 0.2 and weight decay of $10^{-4}$, as prescribed in (Huang et al., 2016). The DenseNet121 and DenseNet201 models have 6.9 million and 18 million parameters, respectively, and are both initialized with He normal (He et al., 2015).
Lastly, we explore a transfer learning scenario on the CIFAR-100 dataset (Krizhevsky, 2012), a choice reflecting the DL community’s growing preference of using pre-trained models in such downstream tasks (Han et al., 2021; Ridnik et al., 2021). For example, a pre-trained visual transformer (ViT) on ImageNet, transferred to classify CIFAR-100, is currently on par with the state-of-the-art results for this task (Dosovitskiy et al., 2020). We adopt this exact transfer learning scenario, leveraging the more powerful ConvNeXtLarge model, pre-trained on ImageNet, with $198$ million parameters (Liu et al., 2022; Chollet et al., 2015). Following the feature extraction step (Goodfellow et al., 2016; Zhuang et al., 2019), the testing accuracy on CIFAR-100 stands at $60\%$. Subsequently, we employ and evaluate our FDA algorithms in the arduous fine-tuning stage, where the entirety of the model is trained (Yosinski et al., 2014).
Algorithms. We consider five distributed deep learning algorithms: LinearFDA, SketchFDA, Synchronous  [^2] The name was derived from the Bulk Synchronous Parallel approach; can be understood as a special case of the FDA Algorithm 1 where $\Theta$ is set to zero., FedAdam (Reddi et al., 2021), and FedAvgM (Hsu et al., 2019); the first three are standard in all experiments. Depending on the local optimizer, Adam (Kingma and Ba, 2017) or SGD with Nesterov momentum (SGD-NM) (Sutskever et al., 2013b), we also include their federated counterparts FedAdam or FedAvgM, respectively.
@@
Hyper-Parameters & Optimizers. Hyper-parameters unique to each training dataset and model are detailed in Table 1; $\Theta$ is pertinent to FDA algorithms and not applicable to others. Notably, a guideline for setting the parameter $\Theta$ is provided in Section 4.2. For experiments involving FedAvgM and FedAdam, we use $E=1$ local epochs, following (Reddi et al., 2021). For experiments with LeNet-5 and VGG16*, local optimization employs Adam, using the default settings as per (Kingma and Ba, 2017). In these cases, FedAdam also adheres to the default settings for both local and server optimization (Reddi et al., 2021; Chollet et al., 2015). For DenseNet121 and DenseNet201, local optimization is performed using SGD with Nesterov momentum (SGD-NM), setting the momentum parameter at $0.9$ and learning rate at $0.1$ (Huang et al., 2016). For FedAvgM, local optimization is conducted with default settings (Hsu et al., 2019; Chollet et al., 2015), while server optimization employs SGD with momentum, setting the momentum parameter and learning rate to $0.9$ and $0.316$, respectively (Reddi et al., 2021). Lastly, for the transfer learning experiments, local optimization leverages AdamW (Loshchilov and Hutter, 2017), with the hyper-parameters used for fine-tuning ConvNeXtLarge in the original study (Liu et al., 2022).
@@
Main Findings. The main findings of our experimental analyses are: (1) LinearFDA and SketchFDA outperform the Synchronous, FedAdam and FedAvgM techniques (their use depends on the local optimizer choice) by 1-2 orders of magnitude in communication, while maintaining equivalent model performance. (2) LinearFDA and SketchFDA also significantly outperform the FedAdam and FedAvgM techniques in terms of computation. (3) The performance of LinearFDA and SketchFDA is comparable in most experiments. SketchFDA provides a more accurate estimator of the variance and leads to fewer synchronizations than LinearFDA, but has a larger communication overhead for its local state (a sketch, compared to a number). SketchFDA significantly outperforms LinearFDA at the transfer learning scenario. (4) The FDA variants remain robust at various data heterogeneity settings, maintaining comparable performance to the IID case.
@@ 4.2. Results
FDA balances Communication vs. Computation. DDL algorithms face a fundamental challenge: balancing the competing demands of computation and communication. Frequent communication accelerates convergence and potentially improves model performance, but incurs higher network overhead, an overhead that may be prohibitive when workers communicate through lower speed connections. Conversely, reducing communication saves bandwidth but risks hindering, or even stalling, convergence. Traditional DDL approaches, like Synchronous, require synchronizing model parameters after every learning step, leading to significant communication overhead but facilitating faster convergence (lower computation cost). This is evident in Figures 12,  3, and 4 (where Synchronous appears in the bottom right — low computation, very high communication). Conversely, Federated Optimization (FedOpt) methods (Reddi et al., 2021) are designed to be communication-efficient, reducing communication between devices (workers) at the expense of increased local computation. Indeed, as shown in Figures 1-4, FedAvgM and FedAdam reduce communication by orders of magnitude but at the price of a corresponding increase in computation. Our two proposed FDA strategies achieve the best of both worlds: the low computation cost of traditional methods and the communication efficiency of FedOpt approaches, as seen in Figures 1-4. In fact, they significantly outperform FedAvgM and FedAdam in their element, that is, communication-efficiency. Across all experiments, the FDA methods’ distributions lie in the desired bottom left quadrant — low computation, very low communication.
@@
FDA counters diminishing returns. The phenomenon of diminishing returns states that as a DL model nears its learning limits for a given dataset and architecture, each additional increment in accuracy may necessitate a disproportionate increase in training time, tuning, and resources (Goodfellow et al., 2016; Thompson et al., 2021). We first clearly notice this with VGG16* on MNIST in Figure 2 for all three data heterogeneity settings. For a $0.001$ increase in accuracy (effectively 10 misclassified testing images) FedAdam needs an order of magnitude more computation and communication. Similarly, Synchronous requires comparable increases in computation and approximately half an order of magnitude more in communication. On the other hand, the FDA methods suffer a slight (if any) increase in computation and communication for this accuracy enhancement. For DenseNet121 and DenseNet201 on CIFAR-10 (Figures  34), FedAvgM and Synchronous require half an order of magnitude more computation and communication to achieve the final marginal accuracy gains ($0.78$ to $0.81$ for DenseNet121, and $0.78$ to $0.8$ for DenseNet201). In contrast, the FDA methods have almost no increase in communication and comparable increase in computation.
@@
Figure 5. Training accuracy progression with a test accuracy target (horizontal line) of $0.8$ (left), and $0.78$ (right). Dashed and doted lines indicate when LinearFDA and SketchFDA attain the target accuracy, respectively. A smaller final gap between training and target accuracy indicates less overfitting, i.e., better generalization capabilities of the trained model
@@
Dependence on $K$. In distributed computing, scaling up typically results in proportional speed improvements. In DDL, however, scalability is less predictable due to the nuanced interplay of computation and communication costs with convergence, complicating the expected linear speedup. This unpredictability is starkly illustrated with VGG16* on the MNIST dataset across all data heterogeneity settings and all strategies. Specifically, Figure 6 demonstrates (at its upper part) that increasing the number of workers does not decrease computation – except for FedAdam which begins with significantly high computation – but rather exacerbates communication (shown at the upper part of the figure). These findings are troubling, as they reveal scaling up only hampers training speed and wastes resources. However, for more complex learning tasks like training DenseNet-121 and DenseNet-201 on CIFAR-10 (upper part of Figures 78), the expected behavior starts to emerge. Specifically, for DenseNet-121, scaling up ($K$ increase) leads to a decrease in computation cost for all strategies. Communication cost, however, increases with $K$ for all methods except Synchronous, which maintains constant communication irrespective of worker count, but at the expense of orders of magnitude higher communication overhead. Notably, while our findings might, in some cases, suggest potential speed benefits of not scaling up (smaller $K$), DDL is increasingly conducted within federated settings, where there is no other choice but to utilize the high number of workers.
@@
FDA: Dependence on $\Theta$. The variance threshold $\Theta$ can be seen as a lever in balancing communication and computation; essentially, it calibrates the trade-off between these two costs. A higher $\Theta$ allows for greater model divergence before synchronization, reducing communication at the cost of potentially increased computation to achieve convergence. This impact of $\Theta$ is consistently observed across all two FDA strategies, learning tasks, and data heterogeneity settings (Figures 67 and 8). Interestingly, for more complex models like DenseNet121 and DenseNet201 on CIFAR-10, increasing the variance threshold ($\Theta$) does not lead to a significant rise in computation cost, as illustrated in Figures 7 and 8. It suggests that the FDA methods, by strategically timing synchronizations (monitoring the variance), substantially reduce the number of necessary synchronizations without a proportional increase in computation for the same model performance; this is particularly promising for complex DDL tasks.
Figure 6. VGG16* on MNIST — Accuracy Target: $0.994$
Figure 7. DenseNet121 on CIFAR-10 — Accuracy Target: $0.8$
Figure 8. DenseNet201 on CIFAR-10 — Accuracy Target: $0.78$
@@
FDA: Choice of $\Theta$. The experimental results suggest that selecting any $\Theta$ within a specific order of magnitude (e.g., between $10^{2}$ and $10^{3}$ for DenseNet201) ensures convergence, as demonstrated in Figures 67, and 8. Therefore, identifying this range becomes crucial. To this end, we conducted extensive exploratory testing to estimate the $\Theta$ ranges for each learning task which are predominantly influenced by the number of parameters $d$ of the DNN. Within this context, $\Theta$ values outside the desirable range exhibit notable effects: below this range, the training process mimics Synchronous or Local-SGD approaches with small $\tau$, while exceeding it leads to non-convergence. Subsequently, having identified the optimal ranges for $\Theta$, we selected diverse values within them for our experimental evaluation (Table 1), thereby investigating different computation and communication trade-offs. For instance, in the ARIS-HPC environment with an InfiniBand connection (up to 56 Gb/s), experiments show that training wall-time (the total time required for the computation and the communication of the DDL) is predominantly influenced by computation cost, rendering communication concerns negligible. In such contexts, lower $\Theta$ values are favored due to their computational efficiency. On the contrary, in FL settings, where communication typically poses the greater challenge, opting for higher $\Theta$ values proves advantageous; reduction in communication achieved with higher $\Theta$ values will translate in a large reduction in total wall-time.
@@
To assist researchers in selecting an optimal variance threshold, Figure 9 presents empirical estimations for $\Theta$ across three distinct learning settings: FL (assuming a common channel of $0.5$Gbps), Balanced (communication-computation equilibrium), and HPC.
@@
Figure 9. Empirical Estimation of the Variance Threshold
FDA: Linear vs. Sketch. In our main body of experiments, across most learning tasks and data heterogeneity settings, the two proposed FDA methods exhibit comparable performance, as illustrated in Figures 123, and  4. This suggests that the precision of the variance approximation is not critical; occasional ”unnecessary” synchronizations do not significantly impact overall performance. However, in all experiments within the more intricate transfer learning scenario, LinearFDA requires approximately $1.5$ times more communication than SketchFDA to fine-tune the deep ConvNeXtLarge model to equivalent performance levels (Figure 10).
In light of these findings, we conclude the following: for straightforward and less demanding tasks, LinearFDA is the recommended option due to its simplicity and lower complexity per local state computation. On the other hand, for intricate learning tasks and deeper models, SketchFDA becomes the preferred choice, particularly if communication-efficiency is paramount.

### v2

@@ 4.1. Setup
Platform. We employ TensorFlow (Abadi et al., 2016), integrated with Keras (Chollet et al., 2015), as the platform for conducting our experiments. We modified TensorFlow to include our FDA variants and all competitive algorithms. All relevant code, figures, and data of this study are available in https://github.com/miketheologitis/FedL-Sync-FDA.
@@
Datasets & Models. The core experiments involve training Convolutional Neural Networks (CNNs) of varying sizes and complexities on two datasets: MNIST (Deng, 2012) and CIFAR-10 (Krizhevsky, 2012). For the MNIST dataset, we employ LeNet-5 (Lecun et al., 1998), composed of approximately 62 thousand parameters, and a modified version of VGG16 (Simonyan and Zisserman, 2015), denoted as VGG16*, consisting of 2.6 million parameters. VGG16* was specifically adapted for the MNIST dataset, a less demanding learning problem compared to ImageNet (Russakovsky et al., 2015), for which VGG16 was designed. In VGG16*, we omitted the $512$-channel convolutional blocks and downscaled the final two fully connected (FC) layers from $4096$ to $512$ units each. Both models use Glorot uniform initialization (Glorot and Bengio, 2010). For CIFAR-10, we utilize DenseNet121 and DenseNet201 (Huang et al., 2017), as implemented in Keras (Chollet et al., 2015), with the addition of dropout regularization layers at rate 0.2 and weight decay of $10^{-4}$, as prescribed in (Huang et al., 2017). The DenseNet121 and DenseNet201 models have 6.9 million and 18 million parameters, respectively, and are both initialized with He normal (He et al., 2015).
Lastly, we explore a transfer learning scenario on the CIFAR-100 dataset (Krizhevsky, 2012), a choice reflecting the DL community’s growing preference of using pre-trained models in such downstream tasks (Han et al., 2021). For example, a pre-trained visual transformer (ViT) on ImageNet, transferred to classify CIFAR-100, is currently on par with the state-of-the-art results for this task (Dosovitskiy et al., 2021). We adopt this exact transfer learning scenario, leveraging the more powerful ConvNeXtLarge model, pre-trained on ImageNet, with $198$ million parameters (Liu et al., 2022; Chollet et al., 2015). Following the feature extraction step (Goodfellow et al., 2016), the testing accuracy on CIFAR-100 stands at $60\%$. Subsequently, we employ and evaluate our FDA algorithms in the arduous fine-tuning stage, where the entirety of the model is trained (Nakandala and Kumar, 2022).
Algorithms. We consider five distributed deep learning algorithms: LinearFDA, SketchFDA, Synchronous  [^2] The name was derived from the Bulk Synchronous Parallel approach; can be understood as a special case of the FDA Algorithm 1 where $\Theta$ is set to zero., FedAdam (Reddi et al., 2021), and FedAvgM (Hsu et al., 2019); the first three are standard in all experiments. Depending on the local optimizer, Adam (Kingma and Ba, 2015) or SGD with Nesterov momentum (SGD-NM) (Sutskever et al., 2013b), we also include their communication-efficient federated counterparts FedAdam or FedAvgM, respectively.
@@
Hyper-Parameters & Optimizers. Hyper-parameters unique to each training dataset and model are detailed in Table 2; $\Theta$ is pertinent to FDA algorithms and not applicable to others. Notably, a guideline for setting the parameter $\Theta$ is provided in Section 4.3. For experiments involving FedAvgM and FedAdam, we use $E=1$ local epochs, following (Reddi et al., 2021). For experiments with LeNet-5 and VGG16*, local optimization employs Adam, using the default settings as per (Kingma and Ba, 2015). In these cases, FedAdam also adheres to the default settings for both local and server optimization (Reddi et al., 2021; Chollet et al., 2015). For DenseNet121 and DenseNet201, local optimization is performed using SGD with Nesterov momentum (SGD-NM), setting the momentum parameter at $0.9$ and learning rate at $0.1$ (Huang et al., 2017). For FedAvgM, local optimization is conducted with default settings (Hsu et al., 2019; Chollet et al., 2015), while server optimization employs SGD with momentum, setting the momentum parameter and learning rate to $0.9$ and $0.316$, respectively (Reddi et al., 2021). Lastly, for the transfer learning experiments, local optimization leverages AdamW (Loshchilov and Hutter, 2019), with the hyper-parameters used for fine-tuning ConvNeXtLarge in the original study (Liu et al., 2022).
@@
Figure 2. LeNet-5 on MNIST. At Non-IID: Label ”0”, the samples of Label ”0” are assigned to few workers. At Non-IID: 60%, 60% of the dataset is sorted and allocated to workers, causing some workers to receive many samples from the same label
Figure 3. VGG16* on MNIST
4.2. Main Findings
The main findings of our experimental analyses are:
  1. (1)
    LinearFDA and SketchFDA outperform the Synchronous, FedAdam and FedAvgM techniques (their use depends on the local optimizer choice) by 1-2 orders of magnitude in communication, while maintaining equivalent model performance.
  2. (2)
    LinearFDA and SketchFDA also significantly outperform the FedAdam and FedAvgM techniques in terms of computation.
  3. (3)
    The performance of LinearFDA and SketchFDA is comparable in most experiments. SketchFDA provides a more accurate estimator of the variance and leads to fewer synchronizations than LinearFDA, but has a larger communication overhead for its local state (a sketch, compared to two numbers). SketchFDA significantly outperforms LinearFDA at the transfer learning scenario.
  4. (4)
    The FDA variants remain robust at various data heterogeneity settings, maintaining comparable performance to the IID case.
@@ 4.3. Results
Figure 6. Training accuracy progression with a test accuracy target (horizontal line) of $0.8$ (left), and $0.78$ (right). Dashed and doted lines indicate when LinearFDA and SketchFDA attain the target accuracy, respectively. A smaller final gap between training and target accuracy indicates less overfitting, i.e., better generalization capabilities of the trained model.
Figure 7. LeNet-5 on MNIST: Varying the Number of Workers and $\Theta$ — Accuracy Target: $0.98$
Figure 8. VGG16* on MNIST: Varying the Number of Workers and $\Theta$ — Accuracy Target: $0.994$
Figure 9. DenseNet121 on CIFAR-10: Varying the Number of Workers and $\Theta$ — Accuracy Target: $0.8$
Figure 10. DenseNet201 on CIFAR-10: Varying the Number of Workers and $\Theta$ — Accuracy Target: $0.78$
@@
FDA balances Communication vs. Computation. DDL algorithms face a fundamental challenge: balancing the competing demands of computation and communication. Frequent communication accelerates convergence and potentially improves model performance, but incurs higher network overhead, an overhead that may be prohibitive when workers communicate through lower speed connections. Conversely, reducing communication saves bandwidth but risks hindering, or even stalling, convergence. Traditional DDL approaches, like Synchronous, require synchronizing model parameters after every learning step, leading to significant communication overhead but facilitating faster convergence (lower computation cost). This is evident in Figures 234, and 5 (where Synchronous appears in the bottom right — low computation, very high communication). Conversely, Federated Optimization (FedOpt) methods (Reddi et al., 2021) are designed to be communication-efficient, reducing communication between devices (workers) at the expense of increased local computation. Indeed, as shown in Figures 2-5, FedAvgM and FedAdam reduce communication by orders of magnitude but at the price of a corresponding increase in computation. Our two proposed FDA strategies achieve the best of both worlds: the low computation cost of traditional methods and the communication efficiency of FedOpt approaches, as seen in Figures 23,  4, and 5. In fact, they significantly outperform FedAvgM and FedAdam in their element, that is, communication-efficiency. Across all experiments, the FDA methods’ distributions lie in the desired bottom left quadrant — low computation, very low communication.
@@
FDA counters diminishing returns. The phenomenon of diminishing returns states that as a DL model nears its learning limits for a given dataset and architecture, each additional increment in accuracy may necessitate a disproportionate increase in training time, tuning, and resources (Goodfellow et al., 2016; Thompson et al., 2021). We first clearly notice this with VGG16* on MNIST in Figure 3 for all three data heterogeneity settings. For a $0.001$ increase in accuracy (effectively 10 misclassified testing images) FedAdam needs an order of magnitude more computation and communication. Similarly, Synchronous requires comparable increases in computation and approximately half an order of magnitude more in communication. On the other hand, the FDA methods suffer a slight (if any) increase in computation and communication for this accuracy enhancement. For DenseNet121 and DenseNet201 on CIFAR-10 (Figures  4, and 5), FedAvgM and Synchronous require half an order of magnitude more computation and communication to achieve the final marginal accuracy gains ($0.78$ to $0.81$ for DenseNet121, and $0.78$ to $0.8$ for DenseNet201). In contrast, the FDA methods have almost no increase in communication and comparable increase in computation.
@@
Dependence on $K$. In distributed computing, scaling up typically results in proportional speed improvements. In DDL, however, scalability is less predictable due to the nuanced interplay of computation and communication costs with convergence, complicating the expected linear speedup. This unpredictability is starkly illustrated with LeNet-5 and VGG16* on the MNIST dataset across all data heterogeneity settings and all strategies. Figures 7, and 8 (left) demonstrate that increasing the number of workers does not decrease computation – except for FedAdam which begins with significantly high computation – but rather exacerbates communication. These findings are troubling, as they reveal scaling up only hampers training speed and wastes resources. However, for more complex learning tasks like training DenseNet-121 and DenseNet-201 on CIFAR-10 (left part of Figures 910), the expected behavior starts to emerge. Especially for DenseNet-121, scaling up ($K$ increase) leads to a decrease in computation cost for all strategies. Communication cost, however, increases with $K$ for all methods except Synchronous, which maintains constant communication irrespective of worker count, but at the expense of orders of magnitude higher communication overhead. Notably, while our findings might, in some cases, suggest potential speed benefits of not scaling up (smaller $K$), DDL is increasingly conducted within federated settings, where there is no other choice but to utilize the high number of workers. Our FDA variants consistently outperform FedAdam, FedAvgM, and Synchronous in communication efficiency, as demonstrated across all experiments in Figures 5-7. Specifically, they require up to 30 times less communication than FedAdam, 4 times less than FedAvgM, and up to 2.5 orders of magnitude less than Synchronous.
@@
FDA: Dependence on $\Theta$. The variance threshold $\Theta$ can be seen as a lever in balancing communication and computation; essentially, it calibrates the trade-off between these two costs. A higher $\Theta$ allows for greater model divergence before synchronization, reducing communication at the cost of potentially increased computation to achieve convergence. This impact of $\Theta$ is consistently observed across all two FDA strategies, learning tasks, and data heterogeneity settings (Figures 7-10). Interestingly, for more complex models like DenseNet121 and DenseNet201 on CIFAR-10, increasing the variance threshold ($\Theta$) does not lead to a significant rise in computation cost, as illustrated in Figures 9 and 10. It suggests that the FDA methods, by strategically timing synchronizations (monitoring the variance), substantially reduce the number of necessary synchronizations without a proportional increase in computation for the same model performance; this is particularly promising for complex DDL tasks.
Figure 11. Empirical Estimation of the Variance Threshold
@@
FDA: Choice of $\Theta$. The experimental results suggest that selecting any $\Theta$ within a specific order of magnitude (e.g., between $10^{2}$ and $10^{3}$ for DenseNet201) ensures convergence, as demonstrated in Figures 7,  89, and 10. Therefore, identifying this range becomes crucial. To this end, we conducted extensive exploratory testing to estimate the $\Theta$ ranges for each learning task which are predominantly influenced by the number of parameters $d$ of the DNN. Within this context, $\Theta$ values outside the desirable range exhibit notable effects: below this range, the training process mimics Synchronous or Local-SGD approaches with small $\tau$, while exceeding it leads to non-convergence. Subsequently, having identified the optimal ranges for $\Theta$, we selected diverse values within them for our experimental evaluation (Table 2), thereby investigating different computation and communication trade-offs. For instance, in the ARIS-HPC environment (https://www.hpc.grnet.gr/en/hardware-2/) with an InfiniBand connection (up to 56 Gb/s), experiments show that training wall-time (the total time required for the computation and the communication of the DDL) is predominantly influenced by the computation cost, rendering communication concerns negligible. In such contexts, lower $\Theta$ values are favored due to their computational efficiency. On the contrary, in FL settings, where communication typically poses the greater challenge, opting for higher $\Theta$ values proves advantageous; reduction in communication achieved with higher $\Theta$ values will translate in a large reduction in total wall-time.
@@
To assist researchers in selecting the variance threshold, Figure 11 presents empirical estimations for $\Theta$ across three distinct learning settings: (1) FL (assuming a common channel of $0.5$Gbps), (2) Balanced communication-computation equilibrium setting, and (3) our HPC environment at the ARIS supercomputer.
@@
FDA: Linear vs. Sketch. In our main body of experiments, across most learning tasks and data heterogeneity settings, the two proposed FDA methods exhibit comparable performance, as illustrated in Figures 234, and 5. This suggests that the precision of the variance approximation is not critical. However, in all experiments within the more intricate transfer learning scenario, LinearFDA requires approximately $1.5$ times more communication than SketchFDA to fine-tune the deep ConvNeXtLarge model to equivalent performance levels (Figure 12). In light of these findings, we conclude the following: for straightforward and less demanding tasks, LinearFDA is the recommended option due to its simplicity and lower complexity per local state computation. On the other hand, for intricate learning tasks and deeper models, SketchFDA becomes the preferred choice, if communication-efficiency is paramount.